TAC Meeting Postmortem. No major issues at the TAC meeting. DanaBalser reviewed the items discussed during the TAC postmortem. The main issue was how to deal with the SRP balance problems and feedback from adding reviewers to AGN and ETP.
PHT. Four items for 17B.
Directors' CSV File. The problem with duplicate proposals was fixed. There still is an issue with the requested time which is listed as zero. Currently the A, B, C, and N time are added to get the requested time for the report which is the requested time in 17B. Probably best to use the requested time from the proposers. MarkClaussen will see what is done for the VLBA. The GBT reports the time requested from the authors.
Science Program. RickLively thinks the bug in the Science program may be fixed already; he will check.
Public Bit. RickLively updated the code to make this work automatically; if this fails we can revert to running the script as in the past.
PHT-->GB PHT Ping. RickLively will update the code to turn off the ping from the PHT-->GB PHT which is no longer used.
GBT. ToneyMinter reported that he will be late with info due to changes in the Sponsored time and some issues with the GB PHT. He will try to get this done before the Directors' review.
VLBA. MarkClaussen reported that he will try to have this done by Friday.